OpenStudy (anonymous):

Simplify by combining like terms. twenty-two times k minus eighteen times k plus 5 plus 9 times k minus thirteen times k

OpenStudy (anonymous):

22K - 18K +5 +9K + 13K = (22 - 18 + 9 - 13)K + 5 = 0K + 5 = 5